# Age-Normalized HRV Features for Non-Invasive Glucose Prediction: A Pilot Sleep-Aware Machine Learning Study

### v1

@@ Abstract
journal: Biomedical Signal Processing and Control
[…]
[addr1]organization=Department of Computer Science & Engineering, Tezpur University, addressline=Napaam, city=Tezpur, postcode=784 028, country=India
{graphicalabstract}[Uncaptioned image]
{highlights}
A pilot study introduced an age-normalization technique for Heart Rate Variability (HRV) features achieving 25.6% improvement in glucose prediction accuracy compared to conventional methods ($R^{2}$ = 0.161 vs. 0.132, p << 0.01).
A preliminary demonstration of sleep-stage specific HRV analysis for glucose monitoring with REM sleep features showed the strongest predictive capability (r = 0.443, p = 0.004).
A Multi-modal machine learning approach evaluated on 43 subjects with systematic ablation studies confirming age-normalization as critical component for initial performance improvements.
Proof-of-concept feasibility was demonstrated with 84.1% predictions within $\pm 1.5$ mmol/L tolerance and single-lead ECG compatibility. Therefore, larger validation studies are required.
Preliminary statistical validation through 5-fold cross-validation showing initial generalizability (CV = 5.9%).However, its clinical applications require further validation.
@@ 2.8 Implementation Details
3 Results and Discussion
3.1 Model Performance with Age-Normalized HRV
The age-normalized model achieved a 5-fold cross-validated coefficient of determination of $R^{2}=0.161\pm 0.010$, with a mean absolute error (MAE) of $0.182\pm 0.006$ and a Pearson correlation of $r=0.409$ ($p<0.001$). Compared to the non-normalized baseline ($R^{2}=0.132$, MAE = 0.185), this represents a significant improvement of +25.6% in $R^{2}$ (+0.034 absolute), as confirmed by the paired $t$-test ($p<0.01$). These preliminary results suggest that age normalization may enhance the predictive utility of heart rate variability (HRV) features for glucose estimation. Further validation in larger cohorts is required to confirm these findings.
3.2 Feature Importance and Ablation Analysis
The top three predictive features were age-normalized HRV metrics from sleep stages, as described in Table 1: hrv_rem_mean_rr_age_normalized ($r=0.443$, $p=0.004$), hrv_ds_mean_rr_age_normalized ($r=0.438$, $p=0.005$), and diastolic blood pressure (DBP, $r=0.437$). The dominance of age-normalized features provides direct evidence for the efficacy of our normalization strategy in mitigating age-related confounding factors.
Rank Feature Correlation ($r$) $p$-value
1 hrv_rem_mean_rr_age_normalized 0.443 0.004
2 hrv_ds_mean_rr_age_normalized 0.438 0.005
3 DBP (mmHg) 0.437 0.005
4 age 0.433 0.005
5 psqi_age 0.433 0.005
3.3 Systematic Ablation Results
Systematic ablation revealed critical architectural contributions (Table 2):
  • 1.
    Age normalization provides +25.6% performance improvement.
  • 2.
  • 3.
  • 4.
    Clinical features alone show negative predictive capability.
Table 2: Ablation study results (15-feature baseline).
Configuration $R^{2}$ MAE Features $\Delta R^{2}$
Full Model 0.161 0.183 15
No Age Normalization 0.132 0.185 15 -0.034
No Sleep HRV 0.108 0.189 15 -0.058
ECG Only 0.075 0.184 15 -0.091
Clinical Only -0.082 0.207 12 -0.248
3.4 Clinical Accuracy and Model Stability
3.5 Sleep-Stage Specific Analysis
Sleep Stage Mean Correlation Standard Deviation Range
REM Sleep 0.321 ± 0.089 High variability 0.156–0.443
Deep Sleep 0.298 ± 0.076 Moderate variability 0.187–0.438
Rapid Sleep 0.267 ± 0.082 Moderate variability 0.143–0.416
This pilot study suggests that age-normalized HRV features may enhance non-invasive glucose prediction accuracy, achieving a preliminary 25.6% improvement over conventional approaches. This methodology addresses a fundamental limitation in autonomic function assessment across diverse age populations, providing a preliminary solution for age-related performance degradation noted in previous HRV studies. Our preliminary $R^{2}$ = 0.161 shows initial promise compared to recent ECG-based glucose detection studies achieving 75-78% binary classification accuracy, while providing continuous prediction rather than discrete categorization. A preliminary tolerance analysis showing 84.1% of predictions within $\pm 1.5$ mmol/L suggests potential, but extensive clinical validation is required before any utility assessment. The superior performance of REM sleep features aligns with the known physiology. REM sleep exhibits variable autonomic activity reflecting complex metabolic regulation, potentially providing enhanced sensitivity to glucose-related autonomic modulation. Parasympathetic dominance of deep sleep offers complementary metabolic insights.
4 Conclusions

### v2

@@ 1 Introduction
2 Results and Discussion
2.1 Model Performance with Age-Normalized HRV
The age-normalized model achieved a 5-fold cross-validated coefficient of determination of $R^{2}=0.161\pm 0.010$, with a mean absolute error (MAE) of $0.182\pm 0.006$ and a Pearson correlation of $r=0.409$ ($p<0.001$). Compared to the non-normalized baseline ($R^{2}=0.132$, MAE = 0.185), this represents a significant improvement of +25.6% in $R^{2}$ (+0.034 absolute), as confirmed by the paired $t$-test ($p<0.01$). The performance comparison across different machine learning models is illustrated in Fig 1, which demonstrates that BayesianRidge regression achieved optimal glucose prediction performance. These preliminary results suggest that age normalization may enhance the predictive utility of heart rate variability (HRV) features for glucose estimation. Further validation in larger cohorts is required to confirm these findings. Cross-validation analysis revealed robust model stability across all folds, as shown in Fig 2. The coefficient of variation was only 5.9%, confirming consistent generalizability with all folds achieving $R^{2}>0.15$.
2.2 Feature Importance and Ablation Analysis
The top three predictive features were age-normalized HRV metrics from sleep stages, as described in Table 1: hrv_rem_mean_rr_age_normalized ($r=0.443$, $p=0.004$), hrv_ds_mean_rr_age_normalized ($r=0.438$, $p=0.005$), and diastolic blood pressure (DBP, $r=0.437$). The dominance of age-normalized features provides direct evidence for the efficacy of our normalization strategy in mitigating age-related confounding factors. The complete ranking of predictive features is presented in Fig 3, which clearly shows that age-normalized HRV features dominate the top positions with strong statistical significance (12 out of 15 features with p<<0.05).
Rank Feature Correlation ($r$) $p$-value
1 hrv_rem_mean_rr_age_normalized 0.443 0.004
2 hrv_ds_mean_rr_age_normalized 0.438 0.005
3 DBP (mmHg) 0.437 0.005
4 age 0.433 0.005
5 psqi_age 0.433 0.005
2.3 Systematic Ablation Results
The systematic ablation study results are summarized in Fig 4, demonstrating that multi-modal integration significantly outperforms single-modality approaches, with the 15-feature configuration proving optimal for glucose prediction.
Systematic ablation revealed critical architectural contributions as described in Table 2:
Table 2: Ablation study results (15-feature baseline).
Configuration $R^{2}$ MAE Features $\Delta R^{2}$
Full Model 0.161 0.183 15
No Age Normalization 0.132 0.185 15 -0.034
No Sleep HRV 0.108 0.189 15 -0.058
ECG Only 0.075 0.184 15 -0.091
Clinical Only -0.082 0.207 12 -0.248
  • Age normalization provides +25.6% performance improvement.
  • Clinical features alone show negative predictive capability.
These architectural contributions are visually summarized in Fig 5, which illustrates the substantial +25.6% performance improvement achieved through age normalization compared to non-normalized features.
2.4 Clinical Accuracy and Model Stability
2.5 Sleep-Stage Specific Analysis
Sleep Stage Mean Correlation Standard Deviation Range
REM Sleep 0.321 ± 0.089 High variability 0.156–0.443
Deep Sleep 0.298 ± 0.076 Moderate variability 0.187–0.438
Rapid Sleep 0.267 ± 0.082 Moderate variability 0.143–0.416
This pilot study suggests that age-normalized HRV features may enhance non-invasive glucose prediction accuracy, achieving a preliminary 25.6% improvement over conventional approaches. This methodology addresses a fundamental limitation in autonomic function assessment across diverse age populations, providing a preliminary solution for age-related performance degradation noted in previous HRV studies. Our preliminary $R^{2}$ = 0.161 shows initial promise compared to recent ECG-based glucose detection studies achieving 75-78% binary classification accuracy, while providing continuous prediction rather than discrete categorization. A preliminary tolerance analysis showing 84.1% of predictions within $\pm 1.5$ mmol/L suggests potential, but extensive clinical validation is required before any utility assessment. The superior performance of REM sleep features aligns with the known physiology. REM sleep exhibits variable autonomic activity reflecting complex metabolic regulation, potentially providing enhanced sensitivity to glucose-related autonomic modulation as described in Table 3. Parasympathetic dominance of deep sleep offers complementary metabolic insights.
3 Conclusions
@@ 4.8 Implementation Details
[…]
Acknowledgements
Funding